# Post-reduction inference for confidence sets of models

### v1

@@
Heather Battey Heather Battey, Yanbo Tang: Department of Mathematics, Imperial College London, 180 Queen’s Gate, London, SW7 2AZ, UK h.battey@imperial.ac.uk, yanbo.tang@imperial.ac.uk Daniel García Rasines Daniel García Rasines: Colegio Universitario de Estudios Financieros, Calle Almansa 101, Madrid 28040, Spain. daniel.garciarasines@cunef.edu  and  Yanbo Tang
@@
Sparsity in a regression context makes the model itself an object of interest, pointing to a confidence set of models as the appropriate presentation of evidence. A difficulty in areas such as genomics, where the number of candidate variables is vast, arises from the need for preliminary reduction prior to the assessment of models. The present paper considers a resolution using inferential separations fundamental to the Fisherian approach to conditional inference, namely, the sufficiency/co-sufficiency separation, and the ancillary/co-ancillary separation. The advantage of these separations is that no direction for departure from any hypothesised model is needed, avoiding issues that would otherwise arise from using the same data for reduction and for model assessment. In idealised cases with no nuisance parameters, the separations extract all the information in the data, solely for the purpose for which it is useful, without loss or redundancy. The extent to which estimation of nuisance parameters affects the idealised information extraction is illustrated in detail for the normal-theory linear regression model, extending immediately to a log-normal accelerated-life model for time-to-event outcomes. This idealised analysis provides insight into when sample-splitting is likely to perform as well as, or better than, the co-sufficient or ancillary tests, and when it may be unreliable. The considerations involved in extending the detailed implementation to canonical exponential-family and more general regression models are briefly discussed. As part of the analysis for the Gaussian model, we introduce a modified version of the refitted cross-validation estimator of Fan et al. (2012), whose distribution theory is exact in an appropriate conditional sense.
@@ 1. Introduction
For regression problems with many potential explanatory variables, Cox (1968) noted that multiple low-dimensional combinations of variables are often compatible with the data. Although an arbitrary choice between well-fitting models is reasonable for prediction, this will typically be misleading in scientific studies in which subject-matter understanding is sought. This perspective, which was emphasised repeatedly over many years (e.g. Cox, 1995; Cox and Snell, 1974, 1989), has acquired renewed importance from a multitude of modern problems, notably in genomics, in which the number of variables for consideration is vast. Routine practice is to report a single model identified by a variable-selection algorithm, the most widely used being the lasso (Tibshirani, 1996). The recommendation of Cox and Battey (2017), by constrast, is to report all models that are compatible with the data at a chosen significance level, i.e. a confidence set of models. Some theoretical considerations associated with this proposal under a high-dimensional regime were developed by Battey and Cox (2018) and Lewis and Battey (2025).
@@
Suppose that a regression model in $p$ variables is correctly specified but over-conditioned, in the sense that the vector of regression coefficients contains many zero entries, whose positions are unknown. In the remainder of the paper, we therefore equate the specification of a model with a set of variable indices and write $\mathcal{E}^{*}\subset[p]$ for the sparse model in $s^{*}<p$ variables that generated the data. When $p$ is not so large as to make the procedure practically infeasible or theoretically unsound, a simple way of constructing confidence sets of models is via a likelihood-ratio test of every low-dimensional subset of variables $\mathcal{E}_{m}\subset[p]:=\{1,\ldots,p\}$ against the encompassing model $[p]$. All models $\mathcal{E}_{m}$, $m=1,\ldots,2^{p}$ not rejected at level $\alpha$ comprise the confidence set of models $\mathcal{M}$, and this, by standard arguments, contains $\mathcal{E}^{*}$ with asymptotic probability $1-\alpha$ in hypothetical repeated use. It is sensible in practice to restrict attention only to those models of reasonable size, substantially reducing the number of assessments to be made.
@@
For the genomics applications we have in mind, $p$ is typically of the order of several thousand, while the number $n$ of individuals is usually in the hundreds or even dozens. The above construction of $\mathcal{M}$ is therefore infeasible without preliminary reduction from $[p]$ to an encompassing model, $\hat{\mathcal{E}}$ say, of more manageable size. The circumflex emphasises that $\hat{\mathcal{E}}$ is constructed from the data; it is not intended as an estimator of $\mathcal{E}^{*}$, a much smaller subset. Any reasonable variable screening procedure can be used for the construction of $\hat{\mathcal{E}}$; see Lewis and Battey (2025) for a theoretical discussion of the implications of that choice.
@@
The key point as far as the present work is concerned is that, since the model $\hat{\mathcal{E}}$ is selected in the light of the data, it fits the observed data better than an arbitrary model of the same size encompassing the one to be tested. For this reason, a likelihood-ratio test of $\mathcal{E}_{m}\subset\hat{\mathcal{E}}$ against $\hat{\mathcal{E}}$ typically rejects too often in hypothetical repeated use, so that the resulting model confidence set $\mathcal{M}$ has a lower coverage probability than suggested by the nominal level of the test. Battey and Cox (2018) used sample splitting as a simple device for guaranteeing the nominal coverage, although it was noted that a potentially better approach would use a minimal sufficiency separation to assess model adequacy (Barndorff-Nielsen and Cox, 1994), bypassing the likelihood ratio test and obviating sample splitting. Essentially the same inferential separation was used by Engen and Lillegård (1997), Lindqvist et al. (2003), Lindqvist and Taraldsen (2005), Lockhart et al. (2007), and Barber and Janson (2022) to assess model fit, but not in the context of the post-reduction inference considered here, where the arguments in favour of inferential separations, relative to a likelihood-ratio or other directed test, are much stronger. The approaches taken in the above references are based on Monte Carlo algorithms to approximately sample from the appropriate subset of the sample space after suitable conditioning. Markov Chain Monte Carlo methods are too computationally expensive for present purposes, as construction of $\mathcal{M}$ entails assessing all low-dimensional subsets of $\hat{\mathcal{E}}$ for their compatibility with the data, a combinatorially large problem.
@@
A key question is over the efficacy of sample splitting relative to the more sophisticated approaches developed here. In simpler contexts, methods powerful against specific alternatives are often valuable in small samples. This raises the possibility that using a large portion of the data for reduction and a small portion for model assessment may be as or more effective than the inferential separations considered here, provided that the test for model assessment specifies a direction for departure from the null, and provided that the sample size for model assessment is sufficiently large that any asymptotic distributions hold to an adequate order of approximation. In the context of inference on the a collection of means, selected on the basis of data from normal distributions, Cox (1975) showed via a theoretical analysis that the efficiency of sample splitting relative to an exact conditional approach is surprisingly high. In the present context, there is a clearer separation of information in the data relevant to the different phases of analysis, so a greater efficiency gain might be expected, although this has to be counterbalanced by the lack of directionality for the full-sample test. These considerations are probed by simulation in §5 for the normal-theory linear regression model.
@@ 2.1. The co-sufficient manifold
$f_{Y}(y;\theta)=\prod_{i=1}^{n}f_{Y_{i}}(y_{i};\theta)=g(s(y);\theta)\prod_{i=1}^{n}h(y_{i})$ (2.1)
@@
$\mathcal{Q}(s^{o})=\{y\in\mathbb{R}^{n}:s(y)=s^{o}\}\subset\mathbb{R}^{n},$ (2.2)
leaving $n-d$ degrees of freedom for variation of $y$ consistent with the constraint $s(y)=s^{o}$, i.e., $\mathcal{Q}(s^{o})$ is a manifold of dimension $n-d$ embedded in $\mathbb{R}^{n}$. Once $s^{o}$ is observed, the notional random variable $Q(s^{o})$ captures the randomness in $Y$ after conditioning on $S=s^{o}$ and is constrained to $\mathcal{Q}(s^{o})$. The implicit general definition of $Q(s^{o})$ is
$\mathop{\mathrm{pr}}(Q(s^{o})\in\mathcal{A})=\mathop{\mathrm{pr}}(Y\in\mathcal{A}|S=s^{o}),$ (2.3)
@@
Barndorff-Nielsen and Cox (1994, p. 29) encourage an understanding of the separation (2.1) for the assessment of model adequacy by hypothetically treating the data as arising in two stages: in the first stage we learn that $S=s^{o}$ which is used for inference on $\theta$. In the second stage we learn that $Y=y^{o}$, which is effectively one draw from the conditional distribution of $Y$ given $S=s^{o}$. If $y^{o}$ is extreme when calibrated against this conditional distribution, that casts doubt on the adequacy of the model. This is the usual analogue of proof by contradiction underpinning the core Fisherian hypothesis tests. In the context of confidence sets of models, the strategy based on falsification of hypotheses does not involve an alternative hypothesis and thereby evades the post-selection inference problem induced through use of the likelihood ratio statistic with a data-dependent encompassing model.
@@ 3.1. Distribution on the co-sufficient manifold
A minimal sufficient statistic when $\sigma^{2}$ is unknown is either $S=(X^{{\rm T}}Y,\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon})$ if the covariate matrix $X$ is treated as fixed, or $S=(\hat{\theta},\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon},X^{{{\rm T}}}X)$ if $X$ is treated as random, where $\hat{\theta}$ is the ordinary least squares estimator of $\theta$ (also the maximum likelihood estimator) and $\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon}$ is the residual sum of squares. Here we treat $X$ as fixed so that $d=d_{\theta}+1$. The information in the data that is not contained in the sufficient statistic can be extracted through orthogonal projection of $Y$ onto $\mathcal{X}^{\perp}$, followed by an additional conditioning on the realised value of $\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon}$. The first conditioning implicit in the projection restricts to the $(n-d_{\theta})$-dimensional linear subspace $\mathcal{X}^{\perp}$, while the second conditioning restricts the lengths of permissible vectors, so that the co-sufficient manifold is a hypersphere embedded in $\mathcal{X}^{\perp}$, which can also be viewed as an $(n-d)$-dimensional sphere embedded in $\mathbb{R}^{n}$. Depending on context, we will refer to the co-sufficient manifold interchangeably as a hypersphere in $\mathcal{X}^{\perp}$ and an $(n-d)$-dimensional sphere in $\mathbb{R}^{n}$, the term hypersphere meaning that there is one dimension missing from the ambient space. Having restricted to $\mathcal{X}^{\perp}$, we can for most purposes think of $\mathcal{Q}(s^{o})$ as being a unit hypersphere in $\mathbb{R}^{n-d_{\theta}}$, except that the coordinates within the larger space $\mathbb{R}^{n}$ are specified by the orientation of $\mathcal{X}^{\perp}$. The distribution on the sphere is uniform under correct specification of the model, as shown by the following calculation, analogous to ones used in the invariant testing literature (e.g. Kariya, 1980; King, 1980).
@@
Consider $W=w(Y)=U^{{\rm T}}Y$, which is normally distributed of mean zero and covariance $U^{{\rm T}}\Sigma U=\sigma^{2}I_{n-d_{\theta}}$. In the transformation $w\mapsto(w/\|w\|,w^{{\rm T}}w)=:(q,r^{2})$, the volume element transforms as
$dw_{1}\cdots dw_{n-d_{\theta}}=\frac{(r^{2})^{\tfrac{n-d_{\theta}}{2}-1}dr^{2}\textstyle{\prod_{j=1}^{n-d_{\theta}-1}}dq_{j}}{2\bigl{(}1-\textstyle{\sum_{j-1}^{n-d_{\theta}-1}}q_{j}^{2}\bigr{)}^{1/2}}=\textstyle{\frac{1}{2}}(r^{2})^{\tfrac{n-d_{\theta}}{2}-1}dr^{2}d\text{Vol}^{n-d}(q),$ (3.2)
@@
$\frac{\Gamma\bigl{(}\frac{n-d_{\theta}}{2}\bigr{)}(q^{{\rm T}}q)^{-(n-d_{\theta})/2}}{2\pi^{(n-d_{\theta})/2}}d\text{Vol}^{n-d}(q),$ (3.3)
@@
This calculation can be generalised to allow for variance components in the form $\Sigma=\sigma^{2}(I_{n}+V(\alpha))$ where $V(\alpha)$ is a known matrix function of a vector parameter $\alpha=(\alpha_{1},\ldots,\alpha_{s})^{{\rm T}}$ with $V(0)=0$. In this case, the appropriate analogue of (3.3) is given by equation (10) of Battey and McCullagh (2024). Alternatively, a transformation to uniformity on the unit hypersphere in $\mathcal{X}^{\perp}$ is $q\mapsto A_{\alpha}^{-1/2}q/(q^{{\rm T}}A_{\alpha}^{-1}q)^{1/2}=:Q^{\prime}$, where $\sigma^{2}A_{\alpha}=U^{{\rm T}}\Sigma U$. There are many tests of uniformity on the sphere, whose properties are well established in the literature, by contrast, tests for compatibility with equation (10) of Battey and McCullagh (2024), if available, are not mainstream.

### v2

@@
Heather Battey Heather Battey, Yanbo Tang: Department of Mathematics, Imperial College London, 180 Queen’s Gate, London, SW7 2AZ, UK h.battey@imperial.ac.uk, yanbo.tang@imperial.ac.uk , Daniel García Rasines Daniel García Rasines: Colegio Universitario de Estudios Financieros, Calle Almansa 101, Madrid 28040, Spain. daniel.garciarasines@cunef.edu and Yanbo Tang
@@
Sparsity in a regression context makes the model itself an object of interest, pointing to a confidence set of models as the appropriate presentation of evidence. A difficulty in areas such as genomics, where the number of candidate variables is vast, arises from the need for preliminary reduction prior to the assessment of models. The present paper considers a resolution using inferential separations fundamental to the Fisherian approach to conditional inference, namely, the sufficiency/co-sufficiency separation, and the ancillary/co-ancillary separation. The advantage of these separations is that no direction for departure from any hypothesised model is needed, avoiding issues that would otherwise arise from using the same data for reduction and for model assessment. In idealised cases with no nuisance parameters, the separations extract all the information in the data solely for the purpose for which it is useful, without loss or redundancy. The extent to which estimation of nuisance parameters affects the idealised information extraction is illustrated in detail for the normal-theory linear regression model, extending immediately to a log-normal accelerated-life model for time-to-event outcomes. This idealised analysis provides insight into when sample-splitting is likely to perform as well as, or better than, the co-sufficient or ancillary tests, and when it may be unreliable. The considerations involved in extending the detailed implementation to canonical exponential-family and more general regression models are briefly discussed. As part of the analysis for the Gaussian model, we introduce a modified version of the refitted cross-validation estimator of Fan et al. (2012), whose distribution theory is tractable in the appropriate conditional sense.
@@ 1. Introduction
For regression problems with many potential explanatory variables, Cox (1968) noted that multiple low-dimensional combinations of variables are often compatible with the data. An arbitrary choice between well-fitting models, while reasonable for prediction, will typically be misleading for scientific understanding, as different models correspond to different scientific explanations. This perspective, which was emphasised repeatedly over many years (e.g. Cox, 1995; Cox and Snell, 1974, 1989), has acquired renewed importance in the light of a multitude of modern areas, notably genomics, in which the number of variables for consideration is vast. Routine practice is to report a single model identified by a variable-selection algorithm, the most widely used being the lasso (Tibshirani, 1996). The recommendation of Cox and Battey (2017), by constrast, was to report all models that are compatible with the data at a chosen significance level, i.e. a confidence set of models. Some theoretical considerations associated with this proposal under a high-dimensional regime were developed by Battey and Cox (2018) and Lewis and Battey (2025). While there are several papers in the epidemiology literature employing the reduction strategy of Cox and Battey (2017), the more important notion of a confidence set of models is yet to gain traction, although some isolated recent papers have pointed to a need to acknowledge model uncertainty (see §2.2). For some epidemiological examples where sparsity-induced uncertainty has been acknowledged in the form of confidence sets of models, see Kartsonaki et al. (2022) and Prosser et al. (2026).
A set of all sparse models compatible with the data, although superficially less appealing than a single one, conveys more honestly the information in the data. In principle, from the sets of compatible models identified, the conclusions might be narrowed down through a carefully designed experiment, or by way of the contradiction argument laid out in a particular context by Battey and Reid (2023, section 5). Cox and Battey (2017) discussed how one might extract compact messages from large confidence sets of models. Specifically, among the multitude of models reported in their supplementary material, two variables, call them $v_{1}$ and $v_{2}$, are present in 96% and 94% of cases. In 78% of the models in which $v_{2}$ is not present, another variable, $v_{3}$, is present in its place, and only 1% of models include neither $v_{2}$ nor $v_{3}$.
@@
Suppose that a regression model in $p$ variables is correctly specified but over-conditioned, in the sense that the vector of regression coefficients contains many zero entries, the positions of the non-zeros being unknown. In the remainder of the paper, we therefore equate the specification of a model with a set of variable indices and write $\mathcal{E}^{*}\subset[p]$ for the sparse model in $s^{*}<p$ variables that generated the data. When $p$ is not so large as to make the procedure practically infeasible or theoretically unsound, a simple way of constructing a confidence set of models is via a likelihood-ratio test of every low-dimensional subset of variables $\mathcal{E}_{m}\subset[p]:=\{1,\ldots,p\}$ against the encompassing model $[p]$. All models $\mathcal{E}_{m}$, $m=1,\ldots,2^{p}$ not rejected at level $\alpha$ comprise the confidence set of models $\mathcal{M}$, and this, by standard arguments, contains $\mathcal{E}^{*}$ with asymptotic probability $1-\alpha$ in hypothetical repeated use. It is sensible in practice to restrict attention only to those models of reasonable size, substantially reducing the number of assessments to be made.
@@
For the genomics applications we have in mind, $p$ is typically of the order of several thousand, while the number $n$ of individuals is usually in the hundreds or even dozens. The above construction of $\mathcal{M}$ is therefore infeasible without preliminary reduction from $[p]$ to an encompassing model, $\hat{\mathcal{E}}$ say, of more manageable size. The circumflex emphasises that $\hat{\mathcal{E}}$ is constructed from the data; it is not intended as an estimator of $\mathcal{E}^{*}$, a much smaller subset. Any reasonable variable screening procedure can be used for the construction of $\hat{\mathcal{E}}$; see Lewis and Battey (2025) for a theoretical discussion of the implications of that choice when the ultimate objective is the construction of $\mathcal{M}$.
The key point as far as the present work is concerned is that, since the comprehensive model $\hat{\mathcal{E}}$ is selected in the light of the data, it fits the observed data better than an arbitrary model of the same size encompassing the one to be tested. For this reason, a likelihood-ratio test of $\mathcal{E}_{m}\subset\hat{\mathcal{E}}$ against $\hat{\mathcal{E}}$ typically rejects too often in hypothetical repeated use, so that the resulting model confidence set $\mathcal{M}$ has a lower coverage probability than suggested by the nominal level of the test. The problem is most pronounced in the context of logistic regression: in high dimensions there are typically many low-dimensional separating hyperplanes among the candidate variables that perfectly separate cases from non-cases in the observed sample. If just one such set was retained through reduction, a logistic regression fitted to $\hat{\mathcal{E}}$ would have perfect fit, and a likelihood ratio test would be unable to detect any $\mathcal{E}^{*}$ that did not also form a separating hyperplane in the same set of data.
@@
Battey and Cox (2018) used sample splitting as a simple device for restoring the nominal coverage, although it was noted that a potentially better approach would use a minimal sufficiency separation to assess model adequacy (Barndorff-Nielsen and Cox, 1994), bypassing the likelihood ratio test and obviating sample splitting. Essentially the same inferential separation was used by Engen and Lillegård (1997), Lindqvist et al. (2003), Lindqvist and Taraldsen (2005), Lockhart et al. (2007), and Barber and Janson (2022) to assess model fit, but not in the context of the post-reduction inference considered here, where the case for such separations over likelihood-ratio or other directed tests is much stronger. The approaches taken in the above references are based on Monte Carlo algorithms to approximately sample from the appropriate subset of the sample space after suitable conditioning. Markov Chain Monte Carlo methods are too computationally expensive for present purposes, as construction of $\mathcal{M}$ entails assessing all low-dimensional subsets of $\hat{\mathcal{E}}$ for their compatibility with the data, a combinatorially large problem.
@@
A key question is over the efficacy of sample splitting relative to the more sophisticated approaches developed here. In simpler contexts, methods powerful against specific alternatives are often valuable in small samples. This raises the possibility that using a large portion of the data for reduction and a small portion for model assessment may be as or more effective than the inferential separations considered here, provided that the test for model assessment specifies a direction for departure from the null, and provided that the sample size for model assessment is sufficiently large that any asymptotic distributions hold to an adequate order of approximation. In the context of inference on the a collection of means, selected on the basis of data from normal distributions, Cox (1975) showed via a theoretical analysis that the efficiency of sample splitting relative to an exact conditional approach is surprisingly high. In the present context, there is a clearer separation of information in the data relevant to the different phases of analysis, so a greater efficiency gain might be expected, although this has to be counterbalanced by the lack of directionality for the full-sample test. These considerations are probed by simulation in §6 for the normal-theory linear regression model. The normal-theory setting is one in which sample splitting has a relatively high efficiency due to the availability of the $F$-test for assessing model adequacy, whose distribution theory is exact and therefore applies at small sample sizes.
2. Related literature and summary of contributions
2.1. Post-selection versus post-reduction inference
The only commonality between the post-reduction inference problem considered here and the more familiar post-selection inference problem of e.g. Lee et al. (2016), Fithian et al. (2017) is that an uncritical re-use of the data leads to erroneous calibration. The most important conceptual difference from the post-selection inference literature is that our inferential goal, the true model, is not data-dependent, as the reduction phase serves only to focus the analysis, and does not dictate the object of inference. Although the analysis is performed under the assumption that $\{\mathcal{E}^{*}\subset\hat{\mathcal{E}}\}$, conditioning on this event barely affects the distribution of test statistics for model assessment, for the latter satisfies, by construction, $\text{pr}(\mathcal{E}^{*}\subset\hat{\mathcal{E}})\approx 1$.
2.2. Model uncertainty
Several other formulations have as their goal the broad objective of reflecting model uncertainty. Most obviously, Bayesian approaches to model selection give rise to Bayesian credible sets of models, whose interpretations are different. In a low-dimensional context, Hansen et al. (2011) proposed something called a model confidence set, but which does not have the usual properties associated with confidence sets. In particular, Hansen et al. (2011) consider the relative explanatory power of pairs of models among an initial candidate set, leading to specification of a reduced set of models with equally good or equally poor fit. By contrast, if no sparse model is compatible with the data, a confidence set of models in the sense of Cox and Battey (2017) would be empty. Lei (2020) proposes an approach based on cross-validation that returns a set of well-fitting candidate models and has a coverage probability guarantee for the model that is optimal in a predictive sense. The definition used by Ferrari and Yang (2015) is the same as ours; since they consider a low-dimensional setting, there are no issues of post-reduction inference.
Breiman (2001) referred to the existence of many alternative predictive models as the Rashomon effect, and used this to challenge the relevance of statistical models, a view that we do not share. See Rudin (2025) for additional perspectives.
A natural proposal for a confidence set of models would involve applying the lasso to bootstrapped samples. This, however, does not have the required coverage properties, as noise variables appear earlier on the lasso path than the last signal variable with high probability (Su et al., 2017; Su, 2018).
2.3. Summary of contributions
The idea of using the sufficiency/co-sufficiency separation for the avoidance of post-reduction miscalibration of confidence sets of regression models was mentioned briefly elsewhere (Battey, 2022; Battey and Cox 2022). The main contribution of the present paper is to make the relevant inferential separations explicit for regression models, to provide geometric insight into the nature of the inferential problem, and to embed into this framework a generalisation of the randomised inference strategy of Rasines and Young (2023). The work also points to more general insights into the foundations of highly parametrised models, which will be developed elsewhere.
@@ 3. Assessment of model adequacy
3.1. Introduction
The discussion to be presented in this section is, at this point, disconnected from the problem outlined in the introduction. The connection will be established in §5 after the presentation of some explicit calculations for the Gaussian linear model in §4.
@@ 3.2. The co-sufficient manifold
(3.1) $f_{Y}(y;\theta)=\prod_{i=1}^{n}f_{Y_{i}}(y_{i};\theta)=g(s(y);\theta)\prod_{i=1}^{n}h(y_{i})$
@@
(3.2) $\mathcal{Q}(s^{o})=\{y\in\mathbb{R}^{n}:s(y)=s^{o}\}\subset\mathbb{R}^{n},$
leaving ndn-d degrees of freedom for variation of $y$ consistent with the constraint $s(y)=s^{o}$, i.e., $\mathcal{Q}(s^{o})$ is a manifold of dimension ndn-d embedded in $\mathbb{R}^{n}$. Once $s^{o}$ is observed, the notional random variable $Q(s^{o})$ captures the randomness in $Y$ after conditioning on $S=s^{o}$ and is constrained to $\mathcal{Q}(s^{o})$. The implicit general definition of $Q(s^{o})$ is
(3.3) $\mathop{\mathrm{pr}}(Q(s^{o})\in\mathcal{A})=\mathop{\mathrm{pr}}(Y\in\mathcal{A}|S=s^{o}),$
@@
Barndorff-Nielsen and Cox (1994, p. 29) encourage an understanding of the separation (3.1) for the assessment of model adequacy by hypothetically treating the data as arising in two stages: in the first stage we learn that $S=s^{o}$ which is used for inference on $\theta$. In the second stage we learn that $Y=y^{o}$, which is effectively one draw from the conditional distribution of $Y$ given $S=s^{o}$. If $y^{o}$ is extreme when calibrated against this conditional distribution, that casts doubt on the adequacy of the model. This is the usual analogue of proof by contradiction underpinning the core Fisherian hypothesis tests. By calibrating $y^{o}$ against the conditional distribution under the model hypothesis of $Y$ given $S=s^{o}$, we avoid the need to compare a smaller model against a larger one, thereby avoiding the post-reduction miscalibration suffered by a likelihood ratio test against a data-dependent encompassing model, as described in §1.
@@ 4.1. Distribution on the co-sufficient manifold
[…]
@@
Consider $W=w(Y)=U^{{\rm T}}Y$, which is normally distributed with mean zero and covariance $U^{{\rm T}}\Sigma U=\sigma^{2}I_{n-d_{\theta}}$ under the postulated model. In the transformation $w\mapsto(w/\|w\|,w^{{\rm T}}w)=:(q,r^{2})$, the volume element transforms as
(4.2) $dw_{1}\cdots dw_{n-d_{\theta}}=\frac{(r^{2})^{\tfrac{n-d_{\theta}}{2}-1}dr^{2}\textstyle{\prod_{j=1}^{n-d_{\theta}-1}}dq_{j}}{2\bigl(1-\textstyle{\sum_{j-1}^{n-d_{\theta}-1}}q_{j}^{2}\bigr)^{1/2}}=\textstyle{\frac{1}{2}}(r^{2})^{\tfrac{n-d_{\theta}}{2}-1}dr^{2}d\text{Vol}^{n-d}(q),$
@@
(4.3) $\frac{\Gamma\bigl(\frac{n-d_{\theta}}{2}\bigr)(q^{{\rm T}}q)^{-(n-d_{\theta})/2}}{2\pi^{(n-d_{\theta})/2}}d\text{Vol}^{n-d}(q),$
@@
This calculation can be generalised to allow for variance components in the form $\Sigma=\sigma^{2}(I_{n}+V(\alpha))$ where $V(\alpha)$ is a known matrix function of a vector parameter $\alpha=(\alpha_{1},\ldots,\alpha_{s})^{{\rm T}}$ with $V(0)=0$. In this case, the appropriate analogue of (4.3) is given by equation (10) of Battey and McCullagh (2024). Alternatively, a transformation to uniformity on the unit hypersphere in $\mathcal{X}^{\perp}$ is $q\mapsto A_{\alpha}^{-1/2}q/(q^{{\rm T}}A_{\alpha}^{-1}q)^{1/2}=:Q^{\prime}$, where $\sigma^{2}A_{\alpha}=U^{{\rm T}}\Sigma U$.